A helium-filled weather balloon has a volume of 773 L at 18.9°C and 758 mmHg. It is released and rises to an altitude of 6.76 km, where the pressure is 388 mmHg and the temperature is –23.1°C.
The volume of the balloon at this altitude is__________ L.
1
Expert's answer
2020-05-01T14:30:59-0400

we have to apply "\dfrac{P_1V_1}{T_1}=\dfrac{P_2V_2}{T_2}"


"\dfrac{773\times758}{291.9}=\dfrac{V_2\times388}{249.9}\\V_2=1292.85 L"